The liquid of density rho in a horizontal pipe wit
Answer #1

from the equation of continuity

A1 v1 = A2 v2

a^2 v1 = b^2 v2

v2 = a^2 v1/ b^2

from the Bernoulis equation

p1 + 1/2 rho v1^2 = p2 + 1/2 rho v2^2

p2 = p1 + 1/2 rho ( v1^2 - v2^2)